You wish to test the following claim (HaHa) at a significance level of α=0.01α=0.01. For the context of this problem, μd=μ2−μ1μd=μ2-μ1 where the first data set represents a pre-test and the second data set represents a post-test.

      Ho:μd=0
      Ha:μd<0

You believe the population of difference scores is normally distributed, but you do not know the standard deviation. You obtain pre-test and post-test samples for n=48n=48 subjects. The average difference (post - pre) is ¯d=−20.2d¯=-20.2 with a standard deviation of the differences of sd=48.2sd=48.2.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic =

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value =
Answer #1

The statistical software output for this problem is:

Hence,

Test statistic = -2.904

p-value = 0.0028
Answer #2

To calculate the test statistic and p-value for this sample, we will use the one-sample t-test formula:

Test Statistic (t) = (¯d - μd) / (sd / √n)

where: ¯d = sample mean of the differences (post - pre) μd = hypothesized population mean difference (in this case, 0) sd = standard deviation of the differences n = sample size

Given the following values: ¯d = -20.2 μd = 0 sd = 48.2 n = 48

Let's calculate the test statistic first:

t = (-20.2 - 0) / (48.2 / √48) t = -20.2 / (48.2 / 6.9282) t = -20.2 / 6.9502 t ≈ -2.908

Now, to find the p-value, we need to use the t-distribution table or a statistical software. Since the p-value corresponds to the area under the t-distribution curve to the left of the test statistic, we are interested in finding P(t < -2.908).

Using a t-distribution table or a statistical software, we find that the p-value for t < -2.908 is approximately 0.0024 (rounded to four decimal places).

So, the test statistic for this sample is approximately -2.908, and the p-value is approximately 0.0024.

answered by: Hydra Master